Metallic barium crystallizes in a body-centered cubic lattice, with one Ba atom per lattice point. If the metallic radius of Ba is 217 pm, what is the volume of the unit cell in pm3 and in cm3?

 pm3

 cm3


1
Expert's answer
2021-11-03T08:14:36-0400

"V = (\frac{4 r}{3^{1/3}})^3"

V = 125856976.239 pm3 or "1.26\times 10^8" pm3

V = 0.0125857 cm3 or "1.26 \times 10^{-2}" cm3